celticcat:

A geometric series has first term x and last term x^2. Its sum is x^3. What is its common ratio?

jhonyy9:

If you know that the sequence is geometric, you can choose any one term in the sequence and divide it by the previous term to find the common ratio.

jhonyy9:

Each term of a geometric sequence increases or decreases by a constant factor called the common ratio.

jhonyy9:

a_{n}=a*r^{n-1}.

jhonyy9:

a_{n}=r*a_{n-1} for every integer n >= 1.